OpenStudy (anonymous):

multiply 1/9 (-3/7) ... GIVING MEDALS

OpenStudy (anonymous):

@KailiTooKute @kayleenxcx @ogjb_ @mathmate @Marki @morganchavez @leonardo0430 @Pamela16 @William14 @ilovewolf @geekfromthefutur @FloppyRawrRawr

OpenStudy (anonymous):

-1/21

OpenStudy (anonymous):

http://web2.0calc.com/#

OpenStudy (anonymous):

how did you get that, what steps @Fullerene

OpenStudy (anonymous):

how do I do that? @karsynnnn

OpenStudy (anonymous):

No. Equal denominators is in addition/subtraction. In multiplying fractions, multiply numerators, then denominators, and simplify. (1*-3)/(9*7) = -3/63 = -1/21

OpenStudy (anonymous):

thankkkkk you

OpenStudy (anonymous):

yw